OpenStudy (anonymous):

A pig farmer wants to enclose 600m^2 for a rectangular pig pen. Three of the sides will consist of fencing, while the remaining sides will consist of an existing stone wall. Two sides of the fence will be of length 'x' metres. a) Write the length of the third side in terms of 'x'. b) Show that the total length 'L' in metres of the fence is given by L= (2x) + 600/x c) What values can 'x' take if the fence length is not to exceed 75cm?

OpenStudy (campbell_st):

call the dimensions x and l then the area is A = x*l or in this question 600 = x*l solve for l this should help you solve the question.